OpenStudy (anonymous):

If \[\alpha \neq \beta\] and \[3\alpha ^{2}-h \alpha-b=0\]\[3\beta ^{2}-h \beta-b=0\], then \[\alpha+ \beta\]=? A. -\frac{ b }{ 3 } B. \frac{ b }{ 3 } C. h D. -\frac{ h }{ 3 } E. \frac{ h }{ 3 }

hartnn (hartnn):

do u realize that means alpha and beta are roots of equation 3x^2-hx-b=0 ?

OpenStudy (anonymous):

no

hartnn (hartnn):

oh, if alpha and beta are roots of equation 3x^2-hx-b=0 ? then u can write \(3\alpha ^{2}-h \alpha-b=0 \\3\beta ^{2}-h \beta-b=0\) here thats given, and since converse is also true....u can conclude that alpha and beta are roots of equation 3x^2-hx-b=0 got that ?

OpenStudy (anonymous):

I am especially weak in roots......

OpenStudy (anonymous):

2. if α and β are the roots of the equation x^(2)-4x-3=0, then α^(2)+αβ+β^(2)=? A. -13 B. 5 C. 13 D. 16 E. 19 3. Find the range of values of k such that the equation x^(2)+(k-2)x+1=0 has real roots. A. k=4 B. 0<k<4 C. \[0 \le k \le4\] D. k<0 or k>4 E. \[k \le 0 \ / k \ge 4\]

hartnn (hartnn):

hey u got 1st answer ?

OpenStudy (anonymous):

no.....I just type a few more questions

hartnn (hartnn):

did u get how alpha and beta are roots of equation 3x^2-hx-b=0

OpenStudy (anonymous):

h^(2)+12b ??

hartnn (hartnn):

what is h^(2)+12b i didn't get you.... i just asked you whether u got my explanation....

OpenStudy (anonymous):

errrr...well, I can't get it...

OpenStudy (anonymous):

\[3\alpha ^{2}-h \alpha-b=0---(1)\] \[3\beta ^{2}-h \beta-b=0---(2)\] (2)-(1), \[3\alpha ^{2}-h \alpha+3\beta ^{2} -hb=0---(3)\] may I do it like this?

hartnn (hartnn):

i don't think that would help.....but let me see...

OpenStudy (anonymous):

however, I cannot get the answer...

hartnn (hartnn):

3a^2-3b^2 -h(a-b) = 0 (a-b) [3(a+b)-h]=0

hartnn (hartnn):

yes! thats it

hartnn (hartnn):

did u get what i've done ?

OpenStudy (anonymous):

well.... 3α^(2)−hα+3β^(2)−hb =3[α^(2)+β^(2)]-h(α+β) isn't it?

hartnn (hartnn):

\(3\alpha ^{2}-h \alpha-3\beta ^{2} +hb=0---(3) \\ 3(\alpha^2-\beta)^2-h(\alpha-\beta)=0 \\ 3(\alpha-\beta)(\alpha+\beta)-h(\alpha-\beta)=0 \\ \text{factor out }\alpha -\beta\)

hartnn (hartnn):

you subtracted incorrectly

hartnn (hartnn):

do (2)-(1) again

hartnn (hartnn):

what u get ?

OpenStudy (anonymous):

3[α^(2)-β^(2)]-h(α-β)=0

hartnn (hartnn):

yeah , now see my explanation above and factor out alpha - beta what u get ?

OpenStudy (anonymous):

3(α−β)(α+β)−h(α−β)=0 (α−β)(α+β)(3-h)=0 is it? I am weak in this part..

hartnn (hartnn):

\(3\alpha ^{2}-h \alpha-3\beta ^{2} +hb=0---(3) \\ 3(\alpha^2-\beta)^2-h(\alpha-\beta)=0 \\ 3(\alpha-\beta)(\alpha+\beta)-h(\alpha-\beta)=0 \\ \text{factor out }\alpha -\beta \\ (\alpha -\beta)(3(\alpha+\beta)-h)=0\)

OpenStudy (anonymous):

okay, got it

OpenStudy (anonymous):

how come I got the answer like this: